# Data Quality Issues in Flare Prediction using Machine Learning Models

### v1

@@ 2.4 Science-Quality Flare List Augmentation
We augment the Science-Quality flare list by attaching NOAA Active Region (AR) numbers to individual flares. Location information is essential for flare prediction, as flares arise from localized magnetic activity, and many flare-forecasting studies rely on AR identifiers to link flares to HARP patches and utilize region-specific physical features. However, the NCEI Science-Quality flare summary product does not include NOAA AR numbers. It only provides precise flare locations for the GOES-R series (GOES-16/17/18/19) beginning on 2017-02-09, expressed in four coordinate systems: Stonyhurst/heliographic (longitude, latitude), Carrington coordinates, heliocentric radial coordinates, and helioprojective Cartesian (HPC) $(x,y)$ goes8-15_readme.
For Science-Quality flares that successfully match to a SWPC-FTP or SSW flare, we directly adopt the AR number from the matched operational report. However, approximately 37% of science-quality flares cannot be matched, or are matched to entries associated with an invalid AR number (0). To extend AR coverage for these cases, we use the pinpoint flare locations provided in the GOES-R series together with NOAA Solar Region Summary (SRS) reports.
The SWPC SRS reports, available daily through the public FTP archive, provide detailed information for each numbered active region, including its position in Stonyhurst/heliographic coordinates. For every Science-Quality flare with valid HPC coordinates, we assign an AR number by identifying the nearest active region in the SRS report using the Euclidean distance in HPC coordinates:
$\text{Distance}\;=\;(X_{\text{AR}}-X_{\text{Flare}})^{2}+(Y_{\text{AR}}-Y_{\text{Flare}})^{2}.$
Among the flares occurring during the interval for which precise location information is available, 93% (11,783 events) can be assigned a nearest AR. For those flares that also have a nonzero AR number reported in SWPC-FTP, 70% share the same AR number, and more than 95% differ by $\leq 6$ units. Table LABEL:tab:nearest_ar_distance summarizes the Euclidean distance (in arcseconds) between flare and nearest AR. To ensure reliability, we retain only AR assignments with distance <250<250 arcsec, increasing the number of valid AR labels for approximately 9% of all flares in this period.
For flares occurring before the availability of GOES-R pinpoint locations, AR numbers are assigned only when a valid match exists in either SWPC-FTP or SSW. Eventually, we assign valid AR numbers to 72% of all Science-Quality flares. This augmented science-quality flare list is used in the model-training comparisons presented in Section 4. A complete data-processing pipeline is provided in the Appendix.
@@ 3.1 Imaging Predictors: HMI/SDO, AIA/SDO, and SOHO/MDI
The imaging dataset of interest is the imaging data provided by the Atmospheric Imaging Assembly (AIA) (lemen2012atmospheric) and Helioseismic and Magnetic Imager (HMI) (scherrer2012helioseismic) instruments from the Solar Dynamics Observatory (SDO) since its launch in 2010 (pesnell2012solar). These instruments capture the full-disk images of the Sun in high spatial resolution (4096 $\times 4096$) with high temporal frequency as well. However, the data from HMI and AIA do not align up one-to-one as the AIA images have a larger field of view, resulting in different pixel sizes. There has been much work in processing the SDO instrument data to help facilitate its use in machine learning research due to these kinds of discrepancies. For example, galvez2019machine curates a data set by removing poor quality images, rescaling the images to a lower and more manageable resolution (512 $\times 512$), and applied corrections due to machine degradation and exposure issues.
@@
In particular, the data product that we focus on is tailored specifically for machine-learning based flare prediction research [Jin et al., 2024 in prep], which processes the existing HMI HARP data with the AIA images (in the 8 channels of 94, 131, 171, 193, 211, 304, 35, and 1600 Ångstroms) to make it the same field of view and coordinates. This data set will contain all B, C, M, and X-class flares from 2010-2024, starting 24 hours before the flare peak-time at a temporal cadence of 12 minutes. The flare peak time is described by the Geostationary Operational Environmental Satellite (GOES) which is operated by the United State’s National Oceanic and Atmospheric Administration (NOAA).
The HMI instrument is used to study oscillations and the magnetic field of the photosphere of the Sun’s surface. The image contain 2D photospheric maps of the 3 orthogonal magnetic components. This is in contrast to the MDI instrument from SOHO, which only recorded 2D Line-of-sight imaging data for the magnetic field, which has been popular for previous machine-learning models. In particular, we are interested in the radial field component of the magnetic image $B_{r}$ channel of the HMI images. This orthogonal component of the magnetic field is not only where some of the summary statistics (like SHARPS that we will discuss further in Section 3.2) for the Sun is derived from, but also the magnetic polarity inversion lines (PIL) can be identified from . This is important as the PIL is known to be one of the most important features in triggering flares and other solar weather activity (sun2023tensor).
One issue regarding the solar imaging data from the HMI and AIA that needs to be worried about is with regards to “near-limb” images that are taken. By “near-limb”, we typically mean the images that are taken of active regions that are on the edge, which is a potential data quality issue. As an example, refer to Table 6, which show the mean proportion of missingness for images near the limb and in the Disk center (which we define to have a longitude of above 60 degrees or less than -60 degrees). As we can see, there is a significantly higher proportion of missings when the images are near the limb, with a lot more variability. Thankfully, we have more images on the disk center that are of “higher-quality”,
@@
For a qualitative perspective of the difference between near limb images and not, Figure 9 is an example of an image of AR11079 of a near limb image. There are clear quality issues with this image, as not only is it blurry and askew, but the specific details and contrast are not very distinct. As such, there is limited information that can possibly be learned by such images.
Figure 9: AR11079 Example Image at 2010-06-13T023100 which is on the limb with a max longitude that is around 80 degrees. It is off-center, blurry, and difficult to make out the specific details.
On the other hand, we have Figure 10 which is of AR11121 when it is near perfectly centered. The quality difference is quite immense, especially as the specific details are quite easy to even visually decipher.
@@
As such, when dealing with these solar flare images, for quality control is best to only deal with the images that are not on the limb. After such a procedure, we can then curate a list of images for the sake of any modeling or other data tasks.
Another concern with the solar flare imaging data is the sporadic missingness we observe in some videos. For many analyses, we would like a continuous 24-hour sequence of an active region’s evolution prior to flare onset. In practice, such complete coverage is uncommon, because the active region must be identified early enough and remain in view for the entire 24 hours. In addition, rare AIA data gaps can occur which mean that even for regions currently being observed, some time intervals are not recorded, creating “gaps” in otherwise continuous video data.
For a visualization to depict the potential negative effects of this data gap, for the task of X hour again prediction of extreme space weather events like solar flares or SEPs, it may be of interest to use historical imaging data of active regions of 24 hours (or 6 hours or anything of interest) before an eruption as training data for a model. However, if we use a flare list listing all flares from 2010-2024 and their associated active region and look at their available images starting from 1 hour away from peak time to 24 hours away, we can see that there is generally less data available the further away we are trying to forecast. This visualization is in Figure 11.
Figure 11: Temporal coverage of on-disk flare images. (Left) Count of unique flares with available data not on the limbs as a function of nearest hour until flare onset. (Right) Stacked annual counts of unique flares that possess complete (non-limb images) 24-hour data coverage, classified by X-ray intensity (M/X, C, and B classes).
@@ 3.2 Vector Predictors: SHARPs
If a machine learning model uses images to predict, then it needs to be trained on a large volume of images in order to predict well. As stated in Section 3.1, HMI images are high-resolution. High-resolution images make for big files and are thus challenging to store in large numbers and to efficiently process in training pipelines. An alternative to using images is to use vectors of summary statistics computed from images; the SHARP dataset is based on this idea. The following description of the dataset draws primarily on (SHARPs).
@@
SHARP stands for Space-weather HMI Active Region Patch. An HMI Active Region Patch, or HARP, is a magnetic structure akin to an active region (HARPs), though a HARP can contain multiple NOAA active regions (SHARPs). A software pipeline automatically detects HARPs in HMI images and tracks them as they traverse the solar disk (HARPs). Every twelve minutes, the pixels within a HARP are used to calculate the vector of sixteen SHARP parameters, each of which describes some aspect of the vector magnetic field of the active region within the HARP. See Table 3 in SHARPs for the definitions and explanations of the SHARPs. For example, the parameter USFLUX measures the total unsigned flux, and the parameter MEANJZD measures the vertical current density. Previous studies have demonstrated that the SHARP parameters have a fair degree of predictive power for forecasting flares (e.g., bobra2015sola; chen2019iden), validating the idea that summary statistics can be an effective alternative to images.
@@
There are two types of HARP products: definitive (DEF-SHARPs) and near-real-time (NRT-SHARPs). See http://jsoc.stanford.edu/HMI/HARPS.html for details. NRT-SHARPs are generated within a few hours of data acquisition and therefore provide more timely access, but their parameter quality is generally lower. Definitive SHARPs become available approximately one week later and incorporate improved calibration and processing. It is important to note that HARP identifiers are not consistent between the definitive and NRT products. The DEF-SHARPs dataset spans an additional two years beyond the availability of NRT-SHARPs (starting on 2010-05-01 versus 2012-09-14), resulting in approximately 2.3 million additional measurement samples that are absent in the NRT set. Even within the overlapping time period, definitive SHARPs include 912 more unique NOAA active regions (ARs) than the NRT product, with only 1043 ARs present in both datasets. Unless otherwise specified, the analyses that follow are based on the definitive SHARPs dataset. We compare these two products in terms of prediction model performance in Section 4.3.
@@
Many other quantities are calculated with the SHARP parameters, such as various longitudes and latitudes. The bits of the QUALITY keyword indicate whether various issues, like eclipses, make the calculated values unreliable; see http://jsoc.stanford.edu/jsocwiki/Lev1qualBits for all the issues that can be flagged. Even if the QUALITY keyword does not indicate any problems, some SHARP parameters can still be missing values; at other times, the entire record, i.e., the SHARP parameters, longitudes and latitudes, QUALITY, etc., may be missing. Keywords that can only be computed when a magnetic field is present can have missing values during “padding intervals”; keywords can also have missing values for HARPs that are faint (HARPs). We do not know if these are the only possible causes of missingness.
@@
We investigated quality issues in the HARP and SHARP data for the period 1 May 2010 to 17 August 2023. The former date is the earliest date for which SHARP data is available (SHARPs). As stated earlier, for any HARP, a new record is created every twelve minutes. The times at which a record can be computed are the start of an hour, 12 minutes after the start, 24 minutes after the start, etc. At each of these times, there may be one or more HARPs in existence, or there may be no HARPs in existence. In the former case, at least one HARP record would be available, and in the latter, no HARP records would be available. For each time in the specified period, we determined whether it lay within the lifetime of some HARP. We found that 526,817 (90%) of the times were within the lifetime of some HARP, and 56,024 (10%) were not within the lifetime of any HARP.
@@
The sequence of times can be partitioned into runs such that within each run, either every time is in some HARP lifetime or no time is in some HARP lifetime. For each run, we computed its start and end times and its duration. These runs are visualized in Figure 12. In the figure, each dot represents one run of times. A run is blue if it consists exclusively of times that lie within the lifetime of some HARP; a run is orange if it only contains times that do not have that property. The first coordinate of a dot is the start time of the corresponding run, while the second coordinate is the run’s duration in days. Orange runs should not be used as there is no HARP data for them and no way to impute it. It seems reasonable to train a model with at least a few months’ worth of data; there are only two blue runs that could be divided into training sets with that much data. Both of those runs correspond to solar maxima.
@@
A HARP record can have one or more problems. A record can be missing values for all SHARP parameters; we call such records missing. A record can be missing values for some, but not all SHARP parameters; we call such records incomplete. We flagged records with these problems. It turned out that all SHARP parameters were missing if and only if the entire HARP record was missing. Another problem is that for certain records, some SHARP parameter is infinite. Yet another is that some records are marked as low-quality. We flagged records with these problems. We also flagged records for which part or all of the HARP was more than 68 degrees from the central meridian. Image quality degrades near the limb of the Sun, so the problems for a record could be due to the HARP being far from the central meridian.
As Table 8 shows, most records don’t have any of the problems described above, but a sizable fraction have some problem. Only 0.1% of records have more than one problem. The most important subsets of problematic records consist of records that are only missing (8.1% of records), records that are only low-quality (5.4% of records), and records that are only incomplete (1.3% of records). Records with an infinite value are virtually nonexistent.
@@
Table 8: A breakdown of the SHARP dataset by data quality problem. A record is missing if the values for all SHARP parameters are missing. A record is incomplete if it has missing values for some SHARP parameters, but not all. A record “has an $\infty$” if the value of some SHARP parameter is $\infty$. A record is low-quality if some bit of the QUALITY keyword is nonzero, which indicates that the record has a quality issue. 85% of records have no problems, and less than 1% have multiple problems. Of those records that do have problems, most are either missing or of low-quality.

### v2

@@ 2.4 Science-Quality Flare List Augmentation
We augment the Science-Quality flare list by attaching NOAA Active Region (AR) numbers to individual flares. Location information is essential for flare prediction, as flares arise from localized magnetic activity, and many flare-forecasting studies rely on AR identifiers to link flares to HARP patches and utilize region-specific physical features. However, the NCEI Science-Quality flare summary product does not include NOAA AR numbers. It only provides precise flare locations for the GOES-R series (GOES-16/17/18/19) beginning on 2017-02-09, expressed in four coordinate systems: Stonyhurst/heliographic (longitude, latitude), Carrington coordinates, heliocentric radial coordinates, and helioprojective Cartesian (HPC) $(x,y)$ Mothersbaugh III et al. (2023).
For Science-Quality flares that are matched to a SWPC-FTP or SSW flare, we directly adopt the AR number from the matched operational report. However, approximately 37% of science-quality flares cannot be matched, or are matched to entries associated with an invalid AR number (0). To extend AR coverage for these cases, we use the pinpoint flare locations provided by the GOES-R series, together with NOAA Solar Region Summary (SRS) reports. The SWPC SRS reports, available daily through the public FTP archive, provide detailed information for each numbered active region, including its position in Stonyhurst/heliographic coordinates. For every Science-Quality flare with valid HPC coordinates, we assign an AR number by identifying the nearest active region in the SRS report using the Euclidean distance in HPC coordinates:
$\text{Distance}\;=\;\sqrt{(X_{\text{AR}}-X_{\text{Flare}})^{2}+(Y_{\text{AR}}-Y_{\text{Flare}})^{2}}.$
Among the flares occurring during the interval for which precise location information is available, 93% (11,783 events) can be assigned to a nearest AR. For those flares that also have a nonzero AR number reported in SWPC-FTP, 70% of the assigned flares share the same AR number with that of the SWPC-FTP, and more than 95% differ by $\leq 6$ units of distance in HPC coordinates. Table 5 summarizes the Euclidean distance (in arcseconds) between a flare and the nearest AR. To ensure reliability, we retain only AR assignments with a distance of <250<250 arcsec. We then assign the nearest AR number to flares that are either unmatched in SWPC-FTP or reported with AR number 0. As a result, 87% of GOES-R flares have valid AR labels, compared with only 77% before augmentation, representing an increase of approximately 10% of all flares in this period.
For flares occurring before the availability of GOES-R pinpoint locations, AR numbers are assigned only when a valid match exists in either SWPC-FTP or SSW. Eventually, we were able to assign valid AR numbers to 72% of all Science-Quality flares. This augmented science-quality flare list is used in the model-training comparisons presented in Section 4. A complete data-processing pipeline is provided in Appendix C.
@@ 3.1 Imaging Predictors: HMI/SDO, AIA/SDO, and SOHO/MDI
The imaging dataset of interest is the imaging data provided by the Atmospheric Imaging Assembly (AIA) (Lemen et al., 2012) and Helioseismic and Magnetic Imager (HMI) (Scherrer et al., 2012) instruments from the Solar Dynamics Observatory (SDO) since its launch in 2010 (Pesnell et al., 2012a). These instruments capture the full-disk images of the Sun in high spatial resolution (4096 $\times 4096$) with high temporal frequency as well. However, the data from HMI and AIA do not align up one-to-one as the AIA images have a larger field of view, resulting in different pixel sizes. There has been extensive work on processing SDO instrument data to facilitate its use in machine learning research, given these discrepancies. For example, Galvez et al. (2019) curates a data set by removing poor quality images, rescaling the images to a lower and more manageable resolution (512 $\times 512$), and applying corrections due to machine degradation and exposure issues.
@@
In particular, [Jin et al., 2026 in prep] processes the existing HMI HARP data with the AIA images (in the 8 channels of 94, 131, 171, 193, 211, 304, 35, and 1600 Ångstroms) to make it the same field of view and coordinates, which is tailored specifically for machine-learning-based flare prediction research. This data set contains all B, C, M, and X-class flares from 2010-2024, starting 24 hours before the flare peak-time at a temporal cadence of 12 minutes. The flare peak time is described by the Geostationary Operational Environmental Satellite (GOES), which is operated by the United States’ National Oceanic and Atmospheric Administration (NOAA). The HMI instrument is used to study oscillations and the magnetic field of the Sun’s photosphere. The image contains 2D photospheric maps of the 3 orthogonal magnetic components. This contrasts with the MDI instrument on SOHO, which recorded only 2D line-of-sight magnetic-field images; HMI contains the radial field component of the magnetic image, $B_{r}$. This orthogonal component of the magnetic field is not only where some of the summary statistics (like the SHARPS that we will discuss further in Section 3.2) for the Sun are derived from, but also where the magnetic polarity inversion lines (PIL) can be identified from. This is important because the PIL is known to be a key driver of flares and other solar weather phenomena (Sun et al., 2023).
One issue with the solar imaging data from the HMI and AIA concerns “near-limb” images, those images of active regions at the limb, which can introduce potential data-quality issues. Table 6 shows the mean proportion of missingness for images near the limb and in the Disk center (which we define to have a longitude of above 60 degrees or less than -60 degrees). As shown, the proportion of missing data is significantly higher when the images are near the limb, with greater variability. Figure 9 is an example near-limb image of AR11079. There are clear quality issues with this image: not only is it blurry and skewed, but the specific details and contrast are not very distinct. In contrast, Figure 10 from AR11121 shows a nearly perfectly centered AR. The quality difference is substantial, especially given that the specific details are readily discernible. As such, when dealing with these solar flare images, it is best to focus on images that are not on the limb for quality control. After such a procedure, we can curate a list of images for use in modeling or other data tasks.
@@
Figure 9: AR11079 Example Image at 2010-06-13T023100 which is on the limb with a max longitude that is around 80 degrees. It is off-center, blurry, and difficult to discern specific details.
@@
Another concern with the solar flare imaging data is the sporadic missingness in some ARs. Ideally, a continuous 24-hour sequence of an active region’s evolution prior to flare onset would be used as the predictors for machine learning models. In practice, such complete coverage is uncommon because the active region must be identified early enough and remain in view for the entire 24 hours. In addition, AIA data gaps can occur, meaning that even for regions currently being observed, some time intervals are not recorded, creating gaps in otherwise continuous video data. Due to the cyclic nature of the Sun’s rotation as well. We will have gaps in observations of active region patches on the Sun because, when they rotate out of our view, we will no longer see them until they rotate back. This means that if there were any solar activity, such as the formation of a new active region or the end of one, the specific time and moment may not be captured.
For the prediction of solar flares, it is desirable to use historical imaging data of active regions 24 hours before an eruption as training data. However, if we use a flare list listing all flares from 2010-2024 and their associated active regions, and examine images available from 24 hours before peak time to peak time, we find that there is generally less data available the further we attempt to forecast. This visualization is in Figure 11.
Figure 11: Temporal coverage of on-disk flare images. (Left) Count of unique flares with available data not on the limbs as a function of the nearest hour until flare onset. (Right) Stacked annual counts of unique flares that possess complete (non-limb images) 24-hour data coverage, classified by X-ray intensity (M/X, C, and B classes).
@@ 3.2 Vector Predictors: SHARPs
If a machine learning model uses images for prediction, it must be trained on a large volume of images to perform well. As stated in Section 3.1, HMI images are high-resolution. High-resolution images generate large files and are thus challenging to store in large quantities and to process efficiently within training pipelines. An alternative to using images is to use vectors of summary statistics computed from images; the SHARP dataset is based on this idea. The following description of the dataset draws primarily on (Bobra et al., 2014).
@@
SHARP stands for Space-weather HMI Active Region Patch. An HMI Active Region Patch, or HARP, is a magnetic structure akin to an active region (JSOC, ), though a HARP can contain multiple NOAA active regions (Bobra et al., 2014). A software pipeline automatically detects HARPs in HMI images and tracks them as they traverse the solar disk (JSOC, ). Every twelve minutes, the pixels within a HARP are used to calculate the vector of sixteen SHARP parameters, each of which describes some aspect of the vector magnetic field of the active region within the HARP. See Table 3 in Bobra et al. (2014) for the definitions and explanations of the SHARPs. For example, the parameter USFLUX measures the total unsigned flux, and the parameter MEANJZD measures the vertical current density. Previous studies have demonstrated that the SHARP parameters exhibit fair predictive power for flare forecasting (e.g., Bobra and Couvidat, 2015; Chen et al., 2019a), supporting the notion that summary statistics can be an effective alternative to images.
@@
There are two types of HARP products: definitive (DEF-SHARPs) and near-real-time (NRT-SHARPs). See http://jsoc.stanford.edu/HMI/HARPS.html for details. NRT-SHARPs are generated within a few hours of data acquisition and therefore provide more timely access, but their parameter quality is generally lower. Definitive SHARPs become available approximately one week later and incorporate improved calibration and processing. It is important to note that HARP identifiers are not consistent between the definitive and NRT products. The DEF-SHARPs dataset spans two additional years beyond the availability of NRT-SHARPs (starting on 2010-05-01 rather than 2012-09-14), yielding approximately 2.3 million additional measurement samples not present in the NRT set. Even within the overlapping period, definitive SHARPs include 912 more unique NOAA active regions (ARs) than the NRT product, with only 1043 ARs present in both datasets. We compare these two products in terms of prediction model performance in Section 4.3.
@@
Many other quantities are computed alongside the SHARP parameters, including various longitudes and latitudes. The bits of the QUALITY keyword indicate whether various issues, like eclipses, make the calculated values unreliable; see http://jsoc.stanford.edu/jsocwiki/Lev1qualBits for all the issues that can be flagged. Even if the QUALITY keyword does not indicate any problems, some SHARP parameters can still be missing; at other times, the entire record, i.e., the SHARP parameters, longitudes and latitudes, QUALITY, etc., may be missing. Keywords that can only be computed when a magnetic field is present can have missing values during “padding intervals”. Keywords can also have missing values for HARPs that are faint (JSOC, ). We do not even know if these are the only possible causes of missingness.
@@
We investigated quality issues in the HARP and SHARP data for the period 1 May 2010 to 17 August 2023. The former date is the earliest date for which SHARP data is available (Bobra et al., 2014). As stated earlier, for any HARP, a new record is created every twelve minutes. The times at which a record can be computed are the start of an hour, 12 minutes after the start, 24 minutes after the start, etc. At each of these times, there may be one or more HARPs in existence, or there may be no HARPs in existence. In the former case, at least one HARP record would be available, and in the latter, no HARP records would be available. For each time in the specified period, we determined whether it lay within the lifetime of some HARP. We found that 526,817 (90%) of the times were within the lifetime of some HARP, and 56,024 (10%) were not within the lifetime of any HARP.
@@
The sequence of times can be partitioned into runs such that, within each run, either every time is in some HARP lifetime or no time is in some HARP lifetime. For each run, we computed its start and end times and its duration. These runs are visualized in Figure 12. In the figure, each dot represents one run of times. A run is blue if it consists exclusively of times that lie within the lifetime of some HARP; a run is orange if it only contains times that do not have that property. The first coordinate of a dot is the start time of the corresponding run, while the second coordinate is the run’s duration in days. Orange runs should not be used, as there is no HARP data for them and no way to impute it. It seems reasonable to train a model with at least a few months’ worth of data; there are only two blue runs that could be divided into training sets with that much data. Both of those runs correspond to solar maxima.
@@
A HARP record can have one or more problems. A record can have missing values for all SHARP parameters and we call such records missing. A record can be missing values for some, but not all SHARP parameters and we call such records incomplete. We flagged records with these problems. It turned out that all SHARP parameters were missing if and only if the entire HARP record was missing. Another problem is that for certain records, some SHARP parameter is infinite. Another is that some records are marked as low quality. We flagged records with these problems. We also flagged records for which part or all of the HARP was more than 68 degrees from the central meridian. Image quality degrades near the limb of the Sun, so the problems for a record could be due to the HARP being far from the central meridian. As Table 8 shows, most records do not have any of the problems described above, but a sizable fraction have some problem. Only 0.1% of records have more than one problem. The most important subsets of problematic records are those that are only missing (8.1% of records), only low-quality (5.4% of records), and only incomplete (1.3% of records). Records with an infinite value are virtually nonexistent.
@@
Table 8: A breakdown of the SHARP dataset by data quality problem. A record is missing if the values for all SHARP parameters are missing. A record is incomplete if it has missing values for some SHARP parameters, but not all. A record “has an $\infty$” if the value of some SHARP parameter is $\infty$. A record is low-quality if some bit of the QUALITY keyword is nonzero, which indicates that the record has a quality issue. 85% of records have no problems, and less than 1% have multiple problems. Of those records that do have problems, most are either missing or of low quality.